# Memory effects in a dynamical decoupling process by parity kicks

### v1

@@ V.3 Numerical simulations
$=$ $\!\text{Tr}_{E}[(\mathbb{I}\otimes e^{\mathcal{L}(t_{b}-t_{a})})(|\psi_{AS}\rangle\langle\psi_{AS}|\otimes\rho_{E}^{\mathrm{th}})]$
where $\mathcal{L}$ is defined in Eq. (54). The Choi-Jamiółkowski matrix $\rho_{\tilde{T}(t_{b},t_{a})}$ is given by
@@
[…]
where $\mathcal{L}_{t^{\prime}}$ is defined in Eq. (57). The time evolutions of the $A\!-\!S\!-\!E$ density matrix are numerically simulated using a fourth-order Runge-Kutta method. In each simulation, the Fock space of the oscillator is truncated to $|n_{\mathrm{max}}\rangle$ where $n_{\mathrm{max}}$ is large enough to guarantee the convergence of all the results, and the kicks are modeled as instantaneous. Moreover, to calculate the strength of the memory effects, the Choi-Jamiółkowski matrix $\rho_{\tilde{T}[n\tau,(n\!-\!1)\tau]\cdots\tilde{T}(2\tau,\tau)\tilde{T}(\tau,0)}$ and $\rho_{\bar{T}[n\tau,(n\!-\!1)\tau]\cdots\bar{T}(2\tau,\tau)\bar{T}(\tau,0)}$ (being the same) are needed. They can be calculated by
@@
The above calculations can be further simplified by $\bar{T}[j\tau,(j\!-\!1)\tau)]=\bar{T}(\tau,0)$ and the property
$\Lambda\rho=d(\text{Tr}\otimes\mathbb{I})[(\rho^{\mathrm{T}}\otimes I)\rho_{\Lambda}]$ (72)
where $\rho^{\mathrm{T}}$ is the transposition of the density matrix $\rho$, $d$ is the dimension of the system and $\rho_{\Lambda}$ is the Choi-Jamiółkowski matrix of a dynamical map $\Lambda$ defined in Sec. III. Now the quantities $\tilde{N}_{M}^{t_{1:n}}$, $\bar{N}_{M}^{t_{1:n}}$, $E$ and $P$ for the dynamical maps can be calculated straightforwardly from the definitions (35), (44), (45) and (51). If we are interested in the final state controlled by the parity kicks when an initial state is prescribed, the strengths of memory effects (47) and (48), the effects of control (48), and the performance (52) can be calculated by numerically simulating the master equation Eq. (54) and (57) and tracing out the environment. Alternatively, they can be calculated directly with the help of Eq. (72) through the Choi-Jamiółkowski matrices Eq. (68)-(V.3) and so on. For example, $\tilde{\rho}_{S}(t_{c})=\tilde{T}(2\tau,0)\rho_{S}(0)=2(\text{Tr}\otimes\mathbb{I})\{[\rho_{S}(0)^{\mathrm{T}}\otimes I]\rho_{\tilde{T}(2\tau,0)}\}$.
@@ V.4 Results for one cycle of parity kicks
In Fig. 2-6 and other further simulations with one cycle of parity kicks, we find that the following properties hold in general (at least) when $t_{c}$ is small enough. (1) The effects of control are close to their upper bounds. (2) The increases of performances are close to the effects of control. (3) The lower bounds of memory effects are close to zero. Therefore, there are $\tilde{N}_{M}^{t_{1:2}}\approx\bar{N}_{M}^{t_{1:2}}\approx 2E\approx 2\Delta P$ and $\tilde{N}_{M}^{t_{1:2}}[\rho_{S}(0)]\approx\bar{N}_{M}^{t_{1:2}}[\rho_{S}(0)]\approx 2E[\rho_{S}(0)]\approx 2\Delta P[\rho_{S}(0)]$. The above properties imply that the increase of the performances (by the control) or the effects of control can directly reflect the strengths of memory effects under certain conditions.
@@ V.5 Results for many cycles of parity kicks
The role of memory effects in more than one cycle of parity kicks ($k\geqslant 2$) is discussed in the following where $\tilde{N}_{M}^{t_{1:n}}$, $\bar{N}_{M}^{t_{1:n}}$, $\tilde{N}_{M}^{t_{1:n}}[\rho_{S}(0)]$ and $\bar{N}_{M}^{t_{1:n}}[\rho_{S}(0)]$ are calculated with $t_{f}=kt_{c}=n\tau$ ($n=2k\geqslant 4$). For simplicity, we will consider scenarios with a large number of cycles and a finite final time $t_{f}$ so that the control is expected to be efficient. Fig. 7 shows the result with $500$ cycles of parity kicks ($t_{f}=500t_{c},n=1000$) where other parameters and settings are the same as those in Fig. 2. It is observed in Fig. 7(a) and (b) that the performances $\Delta P$ and $\Delta P[\rho_{S}(0)]$ under control almost reach their maxima due to the frequent kicks. Meanwhile, the gap between $E^{\mathrm{ub}}$ and $E^{\mathrm{lb}}$ or $E^{\mathrm{ub}}[\rho_{S}(0)]$ and $E^{\mathrm{lb}}[\rho_{S}(0)]$ is much smaller compared with that in Fig. 2(c) or (d), especially when $t_{c}$ is small. The reason is that when the number of cycles is large, the high-frequency kicks efficiently decouple the system from its environment so that the system dynamics is almost unitary and Markovian, i.e., $\tilde{N}_{M}^{t_{1:n}}\approx 0$ and $\tilde{N}_{M}^{t_{1:n}}[\rho_{S}(0)]\approx 0$. Correspondingly, there are $E^{\mathrm{ub}}\approx E^{\mathrm{lb}}\approx\bar{N}_{M}^{t_{1:n}}\approx E\approx\Delta P$ and $E^{\mathrm{ub}}[\rho_{S}(0)]\approx E^{\mathrm{lb}}[\rho_{S}(0)]\approx\bar{N}_{M}^{t_{1:n}}[\rho_{S}(0)]\approx E[\rho_{S}(0)]\approx\Delta P[\rho_{S}(0)]$. For sufficiently frequent kicks, these relations also hold for the regimes in Fig. 3-6 as verified by our simulations (not shown for simplicity). We conclude that in the limit of infinite frequent kicks, the above relations holds exactly. That is, the memory effects without control fully account for the success of the dynamical decoupling task with infinite frequent kicks. In Fig. 7(c) and (d), it is seen that $\bar{N}_{M}^{t_{1:n}}$ or $\bar{N}_{M}^{t_{1:n}}[\rho_{S}(0)]$ may approach its maximum 1 in some cases. Similarly, in a near-Markovian regime like that in Fig. 3, $\bar{N}_{M}^{t_{1:n}}$ or $\bar{N}_{M}^{t_{1:n}}[\rho_{S}(0)]$ might also approach its maximum if the kicks are frequent enough, but not implying strong memory effects. The reason is that longer $t_{f}$ and more initializations of environment (more kicks) are required to obtain a large $\bar{N}_{M}^{t_{1:n}}$ or $\bar{N}_{M}^{t_{1:n}}[\rho_{S}(0)]$ in a near-Markovian regime as discussed before.
@@ VI Conclusions
In this work, we investigate quantitative connections between the environmental memory effects and the characteristics of a dynamical decoupling process, especially, that controlled by parity kicks. Considering that there exists two special time instants (corresponding to two kicks) in one cycle of parity kicks which are linked to the history-past dependence of the system, we tailor our measurements of memory effects to evaluate the strengths of memory effects associated with two specific time instants. For a decoupling process with more than 2 specific time instants (e.g., many cycles of parity kicks), we extend our measurements of memory effects to their multi-time forms. Our measures of memory effects in terms of dynamical maps and quantum states can be interpreted as the history-past dependence of the system dynamics and is easy to calculate. Our main finding is that if the dynamical map of each kick commutes with the system’s dynamical map without control, then the effect of dynamical decoupling ($E$ or $E[\rho_{S}(0)]$) is upper (lower) bounded by the summation (difference) of the strengths of memory effects with and without control. That means (under certain conditions), the environmental memory effects with and without control together determine to what degree the system can be changed, and potentially, the performance (without control) can be increased by the control. On the other hand, the effect of the control or the increase of the performances brought by the control is a manifestation of non-Markovianity.
@@
We apply our theory to a dissipative quantum Rabi model where the qubit is controlled by parity kicks and the oscillator is in a Markovian reservoir at finite temperature. The effects of control, their bounds and the performances with and without control are numerically simulated and compared in different regimes, which verifies our finding. Besides, the relationship among these quantities may take simple forms in particular regimes. For example, in one cycle of parity kicks, there is $\tilde{N}_{M}^{t_{1:2}}\approx\bar{N}_{M}^{t_{1:2}}\approx 2E\approx 2\Delta P$ if the cycle time is short enough. In the case of many cycles, if the frequency of kicks is high enough, there are $\bar{N}_{M}^{t_{1:n}}\approx E\approx\Delta P$ and $\tilde{N}_{M}^{t_{1:n}}\approx 0$. Besides, it is found that the control might be inefficient with particular kick frequencies. Our work provides new insights into the role of memory effects in dynamical decoupling and methods to measure the strength of memory effects by parity kicks.
VII ACKNOWLEDGMENTS

### v2

@@ V.3 Numerical simulations
$=$ $\!\mathrm{Tr}_{E}[(\mathbb{I}\otimes e^{\mathcal{L}(t_{b}-t_{a})})(|\psi_{A\!S}\rangle\langle\psi_{A\!S}|\otimes\rho_{E}^{\mathrm{th}})]$
where $\mathcal{L}$ is defined in Eq. (59) and $|\psi_{AS}\rangle=\frac{1}{\sqrt{2}}(|g\rangle_{A}|g\rangle_{S}+|e\rangle_{A}|e\rangle_{S})$. To calculate $\rho_{\bar{T}(t_{b},t_{a})}$, we first extend Eq. (59) to the ASEA\!-\!S\!-\!E space by replacing $O_{S\!E}$ with $I_{A}\otimes O_{S\!E}$ where $I_{A}$ is the identity operator in the ancillary space and $O_{S\!E}$ represents the operators in Eq. (59). Then, using $\rho_{A\!S\!E}(t_{a})=|\psi_{A\!S}\rangle\langle\psi_{A\!S}|\!\otimes\!\rho_{E}^{\mathrm{th}}$ as the initial state, the extend master equation is solved by a fourth-order Runge-Kutta method from to get $\rho_{A\!S\!E}(t_{b})$. At last, the result is obtained by $\rho_{\bar{T}(t_{b},t_{a})}=\mathrm{Tr}_{E}[\rho_{A\!S\!E}(t_{b})]$. In each simulation, the Fock space of the oscillator is truncated to $|n_{\mathrm{max}}\rangle$ where $n_{\mathrm{max}}$ is large enough to guarantee the convergence of all the results, The Choi-Jamiółkowski matrix $\rho_{\tilde{T}(t_{b},t_{a})}$ is given by
@@
$=$ $\!\mathrm{Tr}_{E}[(\mathbb{I}\!\otimes\!\mathcal{T}\!e^{\int_{t_{a}}^{t_{b}}\!\mathcal{L}_{t^{\prime}}dt^{\prime}})(|\psi_{AS}\rangle\langle\psi_{AS}|\!\otimes\!\rho_{E}^{\mathrm{th}})]$
where $\mathcal{L}_{t^{\prime}}$ is defined in Eq. (62). It is calculated with Eq. (62) with similar steps as done for $\rho_{\bar{T}(t_{b},t_{a})}$ where the kicks are modeled as instantaneous. Moreover, to calculate the strength of the memory effects, the Choi-Jamiółkowski matrix $\rho_{\tilde{T}[n\tau,(n\!-\!1)\tau]\cdots\tilde{T}(2\tau,\tau)\tilde{T}(\tau,0)}$ and $\rho_{\bar{T}[n\tau,(n\!-\!1)\tau]\cdots\bar{T}(2\tau,\tau)\bar{T}(\tau,0)}$ (being the same) are needed. They can be calculated by
@@
The above calculations can be further simplified by $\bar{T}[j\tau,(j\!-\!1)\tau)]=\bar{T}(\tau,0)$ [46] and the property
$\Lambda\rho=d(\mathrm{Tr}\otimes\mathbb{I})[(\rho^{\mathrm{T}}\otimes I)\rho_{\Lambda}]$ (72)
where $\rho^{\mathrm{T}}$ is the transposition of the density matrix $\rho$, $d$ is the dimension of the system and $\rho_{\Lambda}$ is the Choi-Jamiółkowski matrix of a dynamical map $\Lambda$ defined in Sec. III. Now the quantities $\tilde{N}_{M}^{t_{1:n}}$, $\bar{N}_{M}^{t_{1:n}}$, $E$ and $P$ for the dynamical maps can be calculated straightforwardly from the definitions (32), (42), (43) and (49). If we are interested in the final state controlled by the parity kicks when an initial state is prescribed, the strengths of memory effects (45) and (46), the effects of control (46), and the performance (50) can be calculated by numerically simulating the master equation Eq. (59) and (62) and tracing out the environment. Alternatively, they can be calculated directly with the help of Eq. (72) through the Choi-Jamiółkowski matrices Eq. (68)-(V.3) and so on. For example, $\tilde{\rho}_{S}(t_{c})=\tilde{T}(2\tau,0)\rho_{S}(0)=2(\mathrm{Tr}\otimes\mathbb{I})\{[\rho_{S}(0)^{\mathrm{T}}\otimes I]\rho_{\tilde{T}(2\tau,0)}\}$.
V.4 Results for one cycle of parity kicks
@@
In Fig. 1-5 and other further simulations with one cycle of parity kicks, we find that the following properties hold in general (at least) when $t_{c}$ is small enough. (1) The effects of control are close to their upper bounds. (2) The increases of performances are close to the effects of control. (3) The lower bounds of memory effects are close to zero. Therefore, there are $\tilde{N}_{M}^{t_{1:2}}\approx\bar{N}_{M}^{t_{1:2}}\approx 2E\approx 2\Delta P$ and $\tilde{N}_{M}^{t_{1:2}}[\rho_{S}(0)]\approx\bar{N}_{M}^{t_{1:2}}[\rho_{S}(0)]\approx 2E[\rho_{S}(0)]\approx 2\Delta P[\rho_{S}(0)]$. The above properties imply that the increase of the performances (by the control) or the effects of control can directly reflect the strengths of memory effects under certain conditions.
@@ V.5 Results for many cycles of parity kicks
The role of memory effects in more than one cycle of parity kicks ($k\geqslant 2$) is discussed in the following where $\tilde{N}_{M}^{t_{1:n}}$, $\bar{N}_{M}^{t_{1:n}}$, $\tilde{N}_{M}^{t_{1:n}}[\rho_{S}(0)]$ and $\bar{N}_{M}^{t_{1:n}}[\rho_{S}(0)]$ are calculated with $t_{f}=kt_{c}=n\tau$ ($n=2k\geqslant 4$). For simplicity, we will consider scenarios with a large number of cycles and a finite final time $t_{f}$ so that the control is expected to be efficient. Fig. 6 shows the result with $500$ cycles of parity kicks ($t_{f}=500t_{c},n=1000$) where other parameters and settings are the same as those in Fig. 1. It is observed in Fig. 6(a) and (b) that the performances $\Delta P$ and $\Delta P[\rho_{S}(0)]$ under control almost reach their maxima due to the frequent kicks. Meanwhile, the gap between $E^{\mathrm{ub}}$ and $E^{\mathrm{lb}}$ or $E^{\mathrm{ub}}[\rho_{S}(0)]$ and $E^{\mathrm{lb}}[\rho_{S}(0)]$ is much smaller compared with that in Fig. 1(c) or (d), especially when $t_{c}$ is small. The reason is that when the number of cycles is large, the high-frequency kicks efficiently decouple the system from its environment so that the system dynamics is almost unitary and Markovian, i.e., $\tilde{N}_{M}^{t_{1:n}}\approx 0$ and $\tilde{N}_{M}^{t_{1:n}}[\rho_{S}(0)]\approx 0$. Correspondingly, there are $E^{\mathrm{ub}}\approx E^{\mathrm{lb}}\approx\bar{N}_{M}^{t_{1:n}}\approx E\approx\Delta P$ and $E^{\mathrm{ub}}[\rho_{S}(0)]\approx E^{\mathrm{lb}}[\rho_{S}(0)]\approx\bar{N}_{M}^{t_{1:n}}[\rho_{S}(0)]\approx E[\rho_{S}(0)]\approx\Delta P[\rho_{S}(0)]$. For sufficiently frequent kicks, these relations also hold for the regimes in Fig. 2-5 as verified by our simulations (not shown for simplicity). We conclude that in the limit of infinite frequent kicks, the above relations holds exactly. That is, the memory effects without control fully account for the success of the dynamical decoupling task with infinite frequent kicks. In Fig. 6(c) and (d), it is seen that $\bar{N}_{M}^{t_{1:n}}$ or $\bar{N}_{M}^{t_{1:n}}[\rho_{S}(0)]$ may approach its maximum 1 in some cases. Similarly, in a near-Markovian regime like that in Fig. 2, $\bar{N}_{M}^{t_{1:n}}$ or $\bar{N}_{M}^{t_{1:n}}[\rho_{S}(0)]$ might also approach its maximum if the kicks are frequent enough, but not implying strong memory effects. The reason is that longer $t_{f}$ and more initializations of environment (more kicks) are required to obtain a large $\bar{N}_{M}^{t_{1:n}}$ or $\bar{N}_{M}^{t_{1:n}}[\rho_{S}(0)]$ in a near-Markovian regime as discussed before.
@@ VI Conclusion and discussion
In this work, we investigate quantitative connections between the environmental memory effects and the characteristics of a dynamical decoupling process, especially, that controlled by parity kicks. Considering that there exists two special time instants (corresponding to two kicks) in one cycle of parity kicks which are linked to the past-future dependence of the system, we tailor our measurements of memory effects to evaluate the strengths of memory effects associated with two specific time instants. For a decoupling process with more than 2 specific time instants (e.g., many cycles of parity kicks), we extend our measurements of memory effects to their multi-time forms. Our measures of memory effects in terms of dynamical maps and quantum states can be interpreted as the past-future dependence of the system dynamics and is easy to calculate. Our main finding is that if the dynamical map of each kick commutes with the system’s dynamical map without control, then the effect of dynamical decoupling ($E$ or $E[\rho_{S}(0)]$) is upper (lower) bounded by the summation (difference) of the strengths of memory effects with and without control. That means (under certain conditions), the environmental memory effects with and without control together determine to what degree the system can be changed, and potentially, the performance (without control) can be increased by the control. On the other hand, the effect of the control or the increase of the performances brought by the control is a manifestation of non-Markovianity.
@@
We propose a class of Hamiltonians and environment initial states that support the commutation relation for parity kicks. After that, we apply our theory to a dissipative quantum Rabi model where the qubit is controlled by parity kicks and the oscillator is in a Markovian reservoir at finite temperature. The effects of control, their bounds and the performances with and without control are numerically simulated and compared in different regimes, which verifies our finding. Besides, the relationship among these quantities may take simple forms in particular regimes. For example, in one cycle of parity kicks, there is $\tilde{N}_{M}^{t_{1:2}}\approx\bar{N}_{M}^{t_{1:2}}\approx 2E\approx 2\Delta P$ if the cycle time is short enough. In the case of many cycles, if the frequency of kicks is high enough, there are $\bar{N}_{M}^{t_{1:n}}\approx E\approx\Delta P$ and $\tilde{N}_{M}^{t_{1:n}}\approx 0$. Besides, it is found that the control might be inefficient with particular kick frequencies. Our work provides new insights into the role of memory effects in dynamical decoupling and methods to measure the strength of memory effects by parity kicks.
In our model, the Lindblad dissipators in Eq. (59) and Eq. (62) are phenomenological. Microscopically, a standard master equation is derived under the Born-Markov approximation implying a small $\gamma$ compared with $\omega_{E}$ [54]. Besides, the master equations (59) and (62) imply that the dissipation of the harmonic oscillator is independent of the qubit, which is valid when the qubit-oscillator coupling $g$ is not strong [55] compared with $\omega_{S}$, $\omega_{E}$. Therefore, a standard master equation typically can not accurately described the dynamics for large $\gamma$ or $g$. However, our model provides a phenomenological description with highly adjustable non-Markovianity and completely positive dynamics. Besides, such model (and related models) severs as a building block to simulate complex environments as used in the pseudomode theory. Experimentally, Eq. (59) at zero temperature might be effectively achieved by a driven trapped ion system [56].
ACKNOWLEDGMENTS
@@
APPENDIX A DERIVATIONS OF THE BOUND CONDITIONS EQ. (37) AND (41)
In this appendix, we present the derivation of the bound conditions Eq. (37) and (41) for one cycle of parity kicks. Remind that the strength of memory effects with and without one cycle of parity kicks are defined as
$\tilde{N}_{M}^{t_{1:2}}$ $=$ $D[\rho_{\tilde{T}(2\tau,0)},\rho_{\tilde{T}(2\tau,\tau)\tilde{T}(\tau,0)}],$ (73)
$\bar{N}_{M}^{t_{1:2}}$ $=$ $D[\rho_{\bar{T}(2\tau,0)},\rho_{\bar{T}(2\tau,\tau)\bar{T}(\tau,0)}]$ (74)
in Eq. (34) and (35), respectively. Within Eq. (31), it is observed that if $\mathcal{K}_{P}\bar{T}(t_{b},t_{a})=\bar{T}(t_{b},t_{a})\mathcal{K}_{P}$, there is
$\tilde{T}(2\tau,\tau)\tilde{T}(\tau,0)$ $=$ $\mathcal{K}_{P}T(2\tau,\tau)\mathcal{K}_{P}T(\tau,0)$ (75)
$=$ $\mathcal{K}_{P}\mathcal{K}_{P}\bar{T}(2\tau,\tau)\bar{T}(\tau,0)$
$=$ $\bar{T}(2\tau,\tau)\bar{T}(\tau,0)$
where $\mathcal{K}_{P}\mathcal{K}_{P}=\mathbb{I}$ is used due to the properties of parity operators or Eq. (15). Accordingly, the Choi-Jamiółkowski matrices of the left-hand side and right-hand side of Eq. (75) satisfy
$\rho_{\tilde{T}(2\tau,\tau)\tilde{T}(\tau,0)}=\rho_{\bar{T}(2\tau,\tau)\bar{T}(\tau,0)}.$ (76)
Note that the trace distance obeys the triangle inequality
and its variant
$|D(\rho_{1},\rho_{3})-D(\rho_{2},\rho_{3})|\leqslant D(\rho_{1},\rho_{2}).$ (78)
Now let $\rho_{1}=\rho_{\tilde{T}(2\tau,0)}$ and $\rho_{2}=\rho_{\bar{T}(2\tau,0)}$ so that $E=D[\rho_{1},\rho_{2}]=D[\rho_{\tilde{T}(2\tau,0)},\rho_{\bar{T}(2\tau,0)}]$ by the definition of the effect of two parity kicks. If $\mathcal{K}_{P}$ commutes with $\bar{T}(t_{b},t_{a})$, we can set $\rho_{3}$ to be $\rho_{\tilde{T}(2\tau,\tau)\tilde{T}(\tau,0)}$ or $\rho_{\bar{T}(2\tau,\tau)\bar{T}(\tau,0)}$ according to Eq. (76). Specifically, let one of $\rho_{3}$ in Eq. (77) or Eq. (78) be $\rho_{\tilde{T}(2\tau,\tau)\tilde{T}(\tau,0)}$ and the other be $\rho_{\bar{T}(2\tau,\tau)\bar{T}(\tau,0)}$, then Eq. (77) and Eq. (78) together can be expressed as
$|\tilde{N}_{M}^{t_{1:2}}-\bar{N}_{M}^{t_{1:2}}|\leqslant E\leqslant\tilde{N}_{M}^{t_{1:2}}+\bar{N}_{M}^{t_{1:2}}$ (79)
according to the definitions of $\tilde{N}_{M}^{t_{1:2}}$ and $\bar{N}_{M}^{t_{1:2}}$.
For one cycle of parity kicks, the bound condition Eq. (41) in terms of quantum states can be derived similarly. By acting $\tilde{T}(2\tau,0)$, $\tilde{T}(2\tau,\tau)\tilde{T}(\tau,0)$, $\bar{T}(2\tau,0)$ and $\bar{T}(2\tau,\tau)\bar{T}(\tau,0)$ on the initial state $\rho_{S}(0)$, we get 4 final states $\tilde{\rho}_{S}(t_{c})$, $\tilde{\rho}^{\prime}_{S}(t_{c})$, $\bar{\rho}_{S}(t_{c})$ and $\bar{\rho}^{\prime}_{S}(t_{c})$, respectively. If $\mathcal{K}_{P}$ commutes with $\bar{T}(t_{b},t_{a})$, there is
$\tilde{\rho}^{\prime}_{S}(t_{c})=\bar{\rho}^{\prime}_{S}(t_{c}).$ (80)
according to Eq. (75). In Eq. (79), by using the final states $\tilde{\rho}_{S}(t_{c})$, $\tilde{\rho}^{\prime}_{S}(t_{c})$, $\bar{\rho}_{S}(t_{c})$ and $\bar{\rho}^{\prime}_{S}(t_{c})$ instead of $\rho_{\tilde{T}(2\tau,0)}$, $\rho_{\tilde{T}(2\tau,\tau)\tilde{T}(\tau,0)}$, $\rho_{\bar{T}(2\tau,0)}$ and $\rho_{\bar{T}(2\tau,\tau)\bar{T}(\tau,0)}$, respectively, we get the following condition
according to the definitions (39), (40) and $E[\rho_{S}(0)]=D[\tilde{\rho}_{S}(t_{c}),\bar{\rho}_{S}(t_{c})]$.
APPENDIX B DERIVATIONS OF THE BOUND CONDITIONS EQ. (44) AND (47)
In the case of a general dynamical decoupling process with $n$ kicks ($n\geqslant 2$) and $t_{f}=n\tau=kt_{c}$, the bound conditions Eq. (44) and (47) can be derived similarly. Within Eq. (31), if each kick $\mathcal{K}_{j}$ commutes with $\bar{T}(t_{b},t_{a})$, there is
$\tilde{T}[n\tau,(n\!-\!1)\tau]\cdots\tilde{T}(2\tau,\tau)\tilde{T}(\tau,0)$ (82)
$=$ $\mathcal{K}_{n}\bar{T}[n\tau,(\!n\!-\!1\!)\tau]\cdots\mathcal{K}_{2}\bar{T}(2\tau,\tau)\mathcal{K}_{1}\bar{T}(\tau,0)$
$=$ $\mathcal{K}_{n}\cdots\mathcal{K}_{2}\mathcal{K}_{1}\bar{T}[n\tau,(\!n\!-\!1\!)\tau]\cdots\bar{T}(2\tau,\tau)\bar{T}(\tau,0)\quad$
$=$ $\bar{T}[n\tau,(\!n\!-\!1\!)\tau]\cdots\bar{T}(2\tau,\tau)\bar{T}(\tau,0)$
where the properties Eq. (15) is used. It follows that
$\rho_{\tilde{T}[n\tau,(n\!-\!1)\tau]\cdots\tilde{T}(2\tau,\tau)\tilde{T}(\tau,0)}$ (83)
$=$ $\rho_{\bar{T}[n\tau,(\!n\!-\!1\!)\tau]\cdots\bar{T}(2\tau,\tau)\bar{T}(\tau,0)}.$
By using $\rho_{\tilde{T}(n\tau,0)}$, $\rho_{\tilde{T}[n\tau,(n\!-\!1)\tau]\cdots\tilde{T}(2\tau,\tau)\tilde{T}(\tau,0)}$, $\rho_{\bar{T}(n\tau,0)}$ and $\rho_{\bar{T}[n\tau,(\!n\!-\!1\!)\tau]\cdots\!\bar{T}(2\tau,\tau)\bar{T}(\tau,0)}$ instead of $\rho_{\tilde{T}(2\tau,0)}$, $\rho_{\tilde{T}(2\tau,\tau)\tilde{T}(\tau,0)}$, $\rho_{\bar{T}(2\tau,0)}$ and $\rho_{\bar{T}(2\tau,\tau)\bar{T}(\tau,0)}$, respectively in Eq. (79), we obtain
$|\tilde{N}_{M}^{t_{1:n}}-\bar{N}_{M}^{t_{1:n}}|\leqslant E\leqslant\tilde{N}_{M}^{t_{1:n}}+\bar{N}_{M}^{t_{1:n}}$ (84)
according to the definitions (42), (43) and $E=D[\rho_{\tilde{T}(n\tau,0)},\rho_{\bar{T}(n\tau,0)}]$.
Let $\tilde{T}(n\tau,0)$, $\tilde{T}[n\tau,(n\!-\!1)\tau]\cdots\tilde{T}(2\tau,\tau)\tilde{T}(\tau,0)$, $\bar{T}(n\tau,0)$ and $\bar{T}[n\tau,(\!n\!-\!1\!)\tau]\cdots\!\bar{T}(2\tau,\tau)\bar{T}(\tau,0)$ act on the initial state $\rho_{S}(0)$, we get 4 final states $\tilde{\rho}_{S}(kt_{c})$, $\tilde{\rho}^{\prime}_{S}(kt_{c})$, $\bar{\rho}_{S}(kt_{c})$ and $\bar{\rho}^{\prime}_{S}(kt_{c})$, respectively. If each kick $\mathcal{K}_{j}$ commutes with $\bar{T}(t_{b},t_{a})$, there is
$\tilde{\rho}^{\prime}_{S}(kt_{c})=\bar{\rho}^{\prime}_{S}(kt_{c}).$ (85)
according to Eq. (82). Then using $\tilde{\rho}_{S}(kt_{c})$, $\tilde{\rho}^{\prime}_{S}(kt_{c})$, $\bar{\rho}_{S}(kt_{c})$ and $\bar{\rho}^{\prime}_{S}(kt_{c})$ instead of $\rho_{\tilde{T}(n\tau,0)}$, $\rho_{\tilde{T}[n\tau,(n\!-\!1)\tau]\cdots\tilde{T}(2\tau,\tau)\tilde{T}(\tau,0)}$, $\rho_{\bar{T}(n\tau,0)}$ and $\rho_{\bar{T}[n\tau,(\!n\!-\!1\!)\tau]\cdots\!\bar{T}(2\tau,\tau)\bar{T}(\tau,0)}$, respectively in Eq. (84), there is
according to the definitions (42), (43) and $E[\rho_{S}(0)]=D[\tilde{\rho}_{S}(kt_{c}),\bar{\rho}_{S}(kt_{c})]$.